if it takes 6hours to drive 420 kilometer, how long does it take to drive 70 kilometer?
Answer by Sudoku19(4)

Let us say that the amount of speed is the same for the rest then the solution would be:
6x = 420
x = 420 / 6
x = 70
Therefore, x = 70 is considered as 1x = 70.
So the time required it takes to drive a 70 kilometer is just 1 hour, since 1, the coefficient of x, is the amount of hours required to drive a 70 kilometer area.
More so, we could apply ratio.
6:420::x:70
The formula would be 420(x)=6(70)
This would result in 420x = 420
then x = 420 / 420
Thus producing an x (time expressed in unit hours) equal to 1.